Timpco, a retailer, makes both cash and credit sales (i.e., sales on open account). Information regarding budgeted sales for the last quarter of the year is as follows:

October November December
Cash sales $ 150,000 $ 123,000 $ 109,000
Credit sales 150,000 147,600 119,900
Total $ 300,000 $ 270,600 $ 228,900

Past experience shows that 5% of credit sales are uncollectible. Of the credit sales that are collectible, 60% are collected in the month of sale; the remaining 40% are collected in the month following the month of sale. Customers are granted a 1.5% discount for payment within 10 days of billing. Approximately 75% of collectible credit sales take advantage of the cash discount.

Inventory purchases each month are 100% of the cost of the following month’s projected sales. (The gross profit rate for Timpco is approximately 30%.) All merchandise purchases are made on credit, with 20% paid in the month of purchase and the remainder paid in the following month. No cash discounts for early payment are in effect.

Required:

1. Calculate the budgeted total cash receipts for November and December. (Round your intermediate calculations and final answers to the nearest whole dollar amount.)

2. Calculate budgeted cash disbursements for November and December (budgeted total sales for January of the coming year equals $211,000).
Answer #1
Timpco
Budgeted total cash receipts Calculation
Req 1 Nov Dec Oct Nov Dec
Cash sales    123,000    109,000
Collections on account
Oct sales      57,000               -   =150,000*95%*40%
Nov sales      83,186      56,088 =(147,600*95%*60%*75%*98.5%)+(147,600*95%*60%*25%) =147,600*95%*40%
Dec sales               -        67,574 =(119,900*95%*60%*75%*98.5%)+(119,900*95%*60%*25%)
Budgeted total cash receipts    263,186    232,662
Budgeted cash disbursements
Req 2 Oct Nov Dec Oct Nov Dec
Purchases    189,420    160,230    147,700 =270,600*70% '=228,900*70% =211,000*70%
Cash disbursements
Oct Purchases    151,536               -   =189,420*80%
Nov Purchases      32,046    128,184 =160,230*20% =160,230*80%
Dec Purchases               -        29,540 =147,700*20%
Budgeted cash disbursements    183,582    157,724